What is the horizontal force needed to accelerate a 42 kg grocery cart from rest to a velocity of 0.80 m s in 1.75 s


1
Expert's answer
2022-06-09T09:59:30-0400

Given:

"m=42\:\rm kg"

"v_1=0\:\rm m/s"

"v_2=0.80\:\rm m/s"

"t=1.75\:\rm s"



"F=ma=m\frac{v_2-v_1}{t}"

"F=42*\frac{0.80-0}{1.75}=19.2\:\rm N"